OpenStudy (anonymous):

∫01∫0(√1−y2)cos(x2+y2)dxdy evaluate the integral by first converting polar coordinates

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{\sqrt(1-y^2)}\cos(x^2+y^2)dxdy\]

ganeshie8 (ganeshie8):

\(x = r\cos \theta\) \(y = r\sin \theta \) \(dxdy \to rdrd\theta\)

OpenStudy (anonymous):

i found the answer (pi/4)*sin(1)

ganeshie8 (ganeshie8):

whats your double integral after switching to polar ?

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/2}\int\limits_{0}^{1}\cos(r^2)rdrd\]

OpenStudy (anonymous):

i cant open the link :( can you take a photo of it.

ganeshie8 (ganeshie8):

http://prntscr.com/3wzc0w

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)